[figure: 9.7 cm]

CORRELATION BETWEEN NÉEL TEMPERATURE AND SUPEREXCHANGE INTERACTION Fe-O-Fe LINKAGE ANGLES

IN TbFe_{1−x} Al_{x} O_{3}

Andrzej Bombik^{a}, Barbara Leśniewska^{a}, Andrzej W. Pacyna^{b} a Faculty of Phys. Nucl. Techniques, Mining and Metallurgy University, Kraków

b Henryk Niewodniczański Institute of Nuclear Physics, Kraków, Poland

In rare earth orthoferrites REFeO_{3} the Fe^{+3} ions, nearly octahedrally coordinated with oxygen ions, interact through O^{−2} with other Fe^{+3} ions. These superexchange interactions strongly depend on the Fe^{+3}–O^{−2}–Fe^{+3} linkage angle. There are two such linkages: with first and with second neighbours, with corresponding to them θ_{I} and θ_{II} angles (of 140^{o}...150^{o} range). The Fe sublattice Néel temperature is proportional to these angles according to phenomenological formula T_{N} = −\frac{1}{3}T_{N}(θ = 180^{o})(cos θ_{I} +2 cos θ_{II}), where T_{N}(θ = 180^{o}) coefficient is Néel temperature of a hypothetical orthoferrite with linkage angle of 180^{o}. From powder X-ray data, the linkage angles were calculated for the whole TbFe_{1−x} Al_{x} O_{3} series. T_{N} was measured by the differential scanning calorimetry (DSC) for x ≤ 0.4. Both T_{N} and −\frac{1}{3}(cos θ_{I} + 2 cos θ_{II}) quantities are linear functions of x. Comparison of them allows to estimate the value of T_{N}(θ = 180^{o}) ≈752 K.

This work was supported by research grant No 2 P03B 035 14 from State Committee for Scientific Research.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Andrzej W. Pacyna

Address for correspondence : Instytut Fizyki Ja̧drowej im. Henryka Niewodniczańskiego ul. Walerego Eljasza Radzikowskiego 152 31-342 Kraków

Email address : Andrzej.Pacyna@ifj.edu.pl